Parth (parthkohli):

For calculating the identity for \(\rm \large \sin(a + b + c)\), can I use the sum formula twice? Like \(\rm \large \sin((a + b) + c) = \sin(a + b)\cos(c) + \sin(c)\cos(a + b)\) Am I on the right track?

OpenStudy (baldymcgee6):

never heard of such an identity, but maybe they just don't teach it.

Parth (parthkohli):

This is a whole lot complicated than it seems...\[\rm \large (\sin(a)\cos(b)+\sin(b)\cos(a))\cos(c) + \sin(c)(\cos(a)\cos(b)-\sin(a)\sin(b))\]

OpenStudy (baldymcgee6):

just try plugging in values for a, b and c and see if the LHS = RHS

Parth (parthkohli):

lol, I know, just trying to calculate stuff. :)

OpenStudy (baldymcgee6):

We know that Sin(A+B) = SinA CosB + CosA SinB Therefore Sin(A+B+C) = Sin(A+B)CosC + Cos(A+B)SinC Now Sin(A+B) = SinACosB + CosASinC. and Cos(A+B) = CosACosB -SinASinC Substituting we get Sin(A+B+C) = (SinACosB +CosASinB)CosC + (CosACosB -SinASinB)SinC = SinACosBCosC + CosASinBCosC + CosACosBSinC - SinASinBSinC

Parth (parthkohli):

Yes, that's what I am doing... and wow.